OpenStudy (anonymous):

A student planning her curriculum for the upcoming year must select one of seven business courses, one of five mathematics courses, two of eleven elective courses, and either one of four history courses or one of two social science courses. How many different curricula are available for her consideration

OpenStudy (anonymous):

6 * 5 * (see elective odds) * (see history/social odds) odds for the elective is 12C2 12! / 2!(12 - 2)! = 12 * 11 / 2 = 66 odds for one of 2 history or one of 3 social is 5 so the equation... 6 * 5 * 66 * 5 = 9900

OpenStudy (anonymous):

its wrong :(

OpenStudy (anonymous):

11550 is the answer. i got it!